# A Practical Cross-Platform, Multi-Algorithm Study of Quantum Optimisation for Configurational Analysis of Materials

### v2

@@ Abstract
Quantum computers show potential for achieving computational advantage over classical computers, with many candidate applications in the domains of chemistry and materials science. We consider the well-studied problem of configurational analysis of materials and, more specifically, finding the lowest energy configuration of defective graphene structures. This problem acts as a test-case which allows us to study various algorithms that are applicable to Quadratic Unconstrained Binary Optimisation (QUBO) problems, which are generally classically intractable exactly. To solve this problem, we implement two methods, the Variational Quantum Eigensolver (VQE) and Quantum Annealing (QA), on commercially-available gate-based and quantum annealing devices that are accessible via Quantum-Computing-as-a-Service (QCaaS) models. To analyse the performance of these algorithms, we use a toolbox of relevant metrics and compare performance against three classical algorithms. We employ quantum methods to solve fully-connected QUBOs of up to variables, and find that algorithm performance beyond this is restricted by device connectivity, noise and classical computation time overheads. The applicability of our approach goes beyond the selected configurational analysis test-case, and we anticipate that our approach will be of use for optimisation problems in general.
1 Introduction
Quantum mechanics provides the framework for understanding chemistry as we know it. The foundations that let us predict the behaviour and properties of any chemical species have been established for almost a hundred years. That the problem size grows exponentially with the number of particles led Dirac to note, ‘The exact application of these laws leads to equations much too complicated to be soluble’, in 1929 [1]. Hence, despite having a framework, it is difficult for classical computers to simulate chemistry efficiently. Quantum computers offer an alternative platform for solving such problems, building from Feynman’s proposal in 1982 of constructing a computer based on quantum hardware [2].
@@ 1 Introduction
There are several types of chemistry problems that could be considered as candidates to be solved using quantum algorithms, including fermionic simulation [3] and combinatorial optimisation problems. Combinatorial optimisation problems in particular are ubiquitous across science and industry, motivating quantum optimisation as a highly active field of research and development, and a candidate for near-term quantum advantage [4, 5]. In this study we explore the problem of finding the lowest-energy configuration of defective graphene structures, based on the formulation found in Ref. [6]. Suppose we are tasked with removing a specific number of atoms from a graphene sheet of sites. The goal is to find which atoms, when removed, will result in a configuration with the lowest-energy. The configuration of lowest energy will be the one with the maximum number of remaining atom-atom bonds. This is an instance of the densest k-subgraph problem, with many applications beyond configurational analysis [7]. In this work, we consider only the case of removing atoms, and thus .
@@
Motivated by these challenges, in this work we present a systematic testing of performance of VQA and QAA methods on commercially-available quantum hardware: a gate-based QPU (ibm_fez) and a quantum annealer (D-Wave Advantage). We use the graphene defect problem as our test-case. The simplicity of this model and its use in a previous pedagogical study [6] removes a level of complexity in understanding our analysis, making it more accessible. However, most of our analysis applies to optimisation problems in general.
@@
We synthesize an application-focused benchmarking framework for quantum optimisation algorithms. We analyse several performance metrics, carefully selected to ensure the framework applies to both digital and analog methods. While this framework is built around a specific test-case, the selected metrics remain test-case independent. As such, we anticipate the framework will be useful across various problems and research themes.
We apply this benchmarking framework in practice to the selected test-case and at sizes that reveal the limitations in the implementation of VQAs and QAAs.
Briefly, the assumptions and limitations of our work are as follows: we assume a simple energy model, specifically the one studied in Ref. [6], and therefore our work is not directly applicable to solving real-world problems. We encode the problem constraint into the QUBO formalism using penalty terms, rather than solving a hard-constrained version of the problem. Furthermore, we were limited to using some of the best commercially available quantum devices. We did not apply error mitigation or other techniques that could improve algorithm performance (as listed in Section 6).
In Section 2 we present other recent efforts in benchmarking optimisation algorithms and real world implementations. We then formulate our specific problem in Section 3. In Section 4 we describe our methodology for the cross-platform benchmarking including the chosen metrics, whose results we present in Section 5. Finally, in Section 6 we discuss open questions and future directions. Additional details of our methodology are included in the Appendices.
2 Related Work
@@
More specifically, Ref. [6] demonstrates how QAA can be applied to solve a version of the graphene defect use-case formulated as a Densest k-Subgraph (DkS) problem (see Appendix A for more detail on its formulation and complexity). This problem is NP-hard for general and bipartite graphs of degree up to [24]. In [6] an atom lattice is encoded as a fully-connected QUBO and executed on quantum annealers. In our work we execute this on larger lattice sizes, using quantum annealers as well as gate-based quantum computers and classical methods. Another related piece of work [25] studies a slightly different formulation of the problem where the goal is to minimise the number of dangling bonds (i.e. connections between vacancies and carbon atoms in the lattice), again executed on a quantum annealer. The same DkS problem as in [6] has also been investigated in a biology-inspired test-case [26] on randomised bipartite graphs, where the topology of the D-Wave device is identified as the main limitation for applying it on quantum hardware.
@@
Ref. [5] provides a comprehensive review of the current state of quantum optimisation, highlighting the challenges and future prospects of the field. The review outlines best practices for making fair comparisons when solving optimisation problems with different methods. The authors showcase recent publications where optimisation problems with varying numbers of qubits and problem connectivity (known as density) are solved using VQAs, comparing each solution’s quality through the approximation ratio metric. The most common problem connectivity involves 3-regular graphs (where each node is connected to three others), which aligns well with quantum hardware layouts. Studies that explore higher qubit counts typically maintain this low problem connectivity. Increasing connectivity for larger qubit systems requires additional SWAP gates, leading to increased hardware noise. However, Refs. [27, 28, 29, 30] all consider Sherrington-Kirkpatrick problems with 100% density, which can be reformulated as QUBO problems. These studies emphasize both the challenge and the importance of choosing standardised metrics to ensure that benchmarks are fair and results are comparable. The chosen set of metrics must fully capture the diverse behaviors of different algorithms. Our work applies the best practices outlined in this review to a particular problem, which stands out in terms of qubit count and high density compared to the other highlighted studies. We achieve a competitive mean approximation ratio with VQE on a QPU relative to the results reported in Table IV of Ref. [5]. While Ref. [5] primarily focuses on gate-based algorithms, we also address questions regarding how QAAs can be fairly compared to variational gate-based methods.
@@
Other studies [31, 32] compare QAAs and gate-based variational methods for specific use-cases. However, a comprehensive study that includes time comparisons, solution quality, and scaling analysis has yet to be conducted. A recent paper [33] compares the Quantum Approximate Optimisation Algorithm (QAOA) to quantum annealing and claims that QAOA, when run on a real gate-based device, can outperform quantum annealing on a 5-regular max-cut graph problem. On the other hand, Ref. [34] argues that these comparisons could be seen as ambiguous, explaining that the reported results from QAOA were enhanced by post-selection. These studies further highlight the need for transparent and fair benchmarking.
3 Problem Formulation
@@
We are focusing on the problem of finding the minimum energy of defective graphene structures. This problem is discussed in more detail in Appendix A. We model the system as an -site hexagonal structure, as shown visually in Figure 1 ( in this example). The first graphene structure explored is a supercell, a unit cell containing two carbon atoms repeated in a arrangement, which corresponds to an variable QUBO. We also investigate other supercells, which translate to QUBOs of varying size. Each (binary) QUBO variable represents a site containing either a carbon atom or a vacancy.
@@
Each of the carbon atoms in Figure 1 can be removed, creating a vacancy that results in the breaking of carbon-carbon bonds. The breaking of these bonds increases the energy of the system’s ground state configuration. The energy increase depends on where the vacancies are introduced. In particular, and following Ref. [6], the energy of the system depends only on the number of remaining carbon-carbon bonds. In this model we impose periodic boundary conditions, making this structure a -regular graph. The supercell’s 2D geometry is maintained for all configurations.
@@
4 Methodology for Multi-Algorithm Benchmarking
@@ 4.1 Algorithms and Parameters
Many of the methods that we employ here involve hyperparameters whose values must be set. Our hyperparameter search finds the set of parameters for each method that return the highest probability of sampling the ground state. A hyperparameter search could be performed aiming to optimise performance with respect to another metric such as runtime. Our hyperparameter optimisation is a grid-based search, conducted by defining feasible search spaces and checking different combinations of values. The results of the hyperparameter search for all methods in this section are provided in Appendix B.
@@ 4.2 Post-Selection
To perform post-selection for our use-case with a single constraint, a constant number of bitstrings must be checked, equal to the number of samples (shots), which have entries each. Therefore, the scaling of this method is linear in . The time required for post-selection at the QUBO problem size, was found to be approximately for samples, which is the maximum number of samples produced as output in any of our experiments. However, for larger problems, this number of samples may need to scale with to achieve non-zero probability of the optimal solution. As increases, this probability decreases due to increased noise from using more qubits, which we have confirmed experimentally.
@@
4.3 Performance Metrics
Defining good benchmarking strategies for measuring performance of optimisation algorithms has been extensively studied in both quantum and classical computing contexts [51]. In general, these strategies fall under two distinct categories. The first category involves allocating a fixed amount of resources, such as time or energy, and measuring the quality of solution obtained, using a well-defined metric for it. The second involves measuring the amount of a resource, using an appropriate metric, that is necessary to acquire a solution that achieves a pre-specified threshold (or target) for the quality of solution.
@@ 4.3 Performance Metrics
In our case we calculate AR using the solutions after post-selection. The reason we do this is because before post-selection we can have energies outside the range . These and values are calculated using constrained brute-force techniques. The AR places the average energy found within the context of the whole energy landscape of the problem. We note that there is some variation in the name and precise definition used in the literature for this metric. When exploring new large problems, the theoretical minimum and maximum values of energy may not be readily available without having to solve using an exact method. An option for larger problems is that, in place of the AR, the values of the the average energy and the minimum energy found experimentally by the particular method can be used individually as metrics.
@@
Having discussed the selected metrics, we also note a number of other metrics that are used in the wider literature. Ref. [32], considers the number of violated constraints. For our QUBO there is only one constraint specified - the number of vacancies. The number of times it is violated could be reflected in a ‘validity’ metric: if many solutions violate the constraint then the validity returned will be low. We instead choose to encode information about constraint violation by post-selecting the data to remove solutions that do not obey the constraint, as described above. A metric specific to the quality of the quantum annealing process in current devices is the percentage of broken chains. We report this value in Appendix E, Figure 19(d).
Another metric often used is the time-to-solution (TTS), where this is the time taken for a method to find the optimal solution once with a desired accuracy (often ). TTS, as defined in Ref. [56], is:
@@
Here we create our collection of solutions by aggregating shots across many experiments. This means that the total number of solutions for quantum annealing and VQE is equal to , with being the number of times the experiment is repeated and the number of shots per experiment. While in simulated annealing and random sampling it is equal to the number of times we run the algorithm. For the definitions of shots and experiments see Section 4.1. This is necessary with quantum annealing as the D-Wave devices have a maximum QPU access time, and in order to increase the number of shots to a sufficiently large number, one needs to run the experiment again[58].
@@
5 Results
We first solve the variable QUBO and analyse our methods’ performance using the metrics defined in Section 4.3. Figure 3 shows the energy probability distributions obtained by each method before post-selection.

### v4

# Cross-Platform Benchmarking of Near‑Term Quantum Optimisation Algorithms

@@ Abstract
Quantum computers show potential for achieving computational advantage over classical computers, with many candidate applications in combinatorial optimisation. We present an application level benchmarking framework for near-term quantum optimisation algorithms using a dense Quadratic Unconstrained Binary Optimisation (QUBO) materials science problem as a representative test-case. To solve this problem, we implement two methods, the Variational Quantum Eigensolver (VQE) and Quantum Annealing (QA), on commercially-available gate-based and quantum annealing devices that are accessible via Quantum-Computing-as-a-Service (QCaaS) models. To analyse the performance of these algorithms, we use a toolbox of relevant metrics and compare performance against three classical algorithms. We employ quantum methods to solve fully-connected QUBOs of up to variables, and find that algorithm performance beyond this is restricted by device connectivity, noise and classical computation time overheads. The applicability of our approach goes beyond the selected configurational analysis test-case, and we anticipate that our approach will be of use for optimisation problems in general.
I Introduction
Combinatorial optimisation problems are ubiquitous across science and industry, motivating quantum optimisation as a highly active field of research and development, and a candidate for near-term quantum advantage [8, 25]. With respect to quantum advantage and performance guarantees, algorithms can be categorised into three classes: provably exact algorithms, which guarantee the optimal solution; provably approximate algorithms with guaranteed performance bounds (such as Decoded Quantum Interferometry [61]); and heuristic algorithms, which offer no performance guarantees. Near-term quantum optimisation algorithms fall into the last category.
For heuristic quantum algorithms, claims of quantum advantage cannot be established through asymptotic complexity arguments and instead require experimental validation. This motivates the need for rigorous benchmarking, in which state of the art quantum and classical methods are compared in a fair and transparent way. Our work addresses this need through the full disclosure of key metrics and methodologies, an aspect that is currently lacking in the field. Cross-platform benchmarking - where a candidate problem believed to be classically hard with practical relevance is solved with competing methods and devices - provides a natural framework for such comparisons. This not only enables the testing of algorithms, but can help track the progress of the underlying quantum hardware that the algorithms are deployed on. As well as informing improvements to the hardware, cross-platform benchmarking can motivate algorithm development where tweaks can be made to improve performance. Our work contributes to ongoing efforts to assess quantum optimisation methods on practically motivated problem instances [39, 27, 77].
@@ I Introduction
To perform our cross-platform benchmarking we chose the problem of finding the lowest-energy configuration of defective graphene structures, based on the formulation found in Ref. [28]. We use the graphene defect problem solely as a representative, high density instance on which to test our benchmarking framework. Suppose we are tasked with removing a specific number of atoms from a graphene sheet of sites. The goal is to find which atoms, when removed, will result in a configuration with the lowest-energy. The configuration of lowest energy will be the one with the maximum number of remaining atom-atom bonds. This is an instance of the densest k-subgraph problem, with many applications beyond configurational analysis [65]. In this work, we consider only the case of removing atoms, and thus .
@@
Motivated by these challenges, in this work we present a systematic testing of performance of VQA and QAA methods on commercially-available quantum hardware: a gate-based QPU (ibm_fez) and a quantum annealer (D-Wave Advantage 6.4). We use the graphene defect problem as our test-case. The simplicity of this model and its use in a previous pedagogical study [28] removes a level of complexity in understanding our analysis, making it more accessible. However, most of our analysis applies to optimisation problems in general.
@@
We introduce a benchmarking framework for quantum optimisation that enables fair and consistent comparison of gate-based variational algorithms, quantum annealing and classical algorithms on the same QUBO instances.
We apply this benchmarking framework in practice to the graphene defect problem on QCaaS hardware, using this hardware to solve fully connected QUBOs up to 72 variables.
We explore the scaling behaviour of the algorithms, finding an efficient classical solution with simulated annealing that scales polynomially with system size . Using simulated annealing, we solved problems with up to 388 variables. We also present some provisional evidence of quantum annealing scaling in polynomial time on the problem sizes that were feasible to run on hardware.
Our work reveals key technical details such as: how higher probability of obtaining the optimal solution can be obtained by using small penalty coefficients, , which keep the QUBO energy range small, in combination with our post-selection, which removes any solutions that violate the constraint (due to using a small ); the large overhead embedding poses to quantum annealing and how this can be reduced with a more efficient technique; the large standard deviation error in optimal solution probability due to differing random initial parameters for VQE; how noise affects VQE convergence and increases the difficulty of finding good variational parameters; how device connectivity limits performance on larger problems; and how classical computation overheads limit the number of experimental repeats that can be excecuted on the IBM QPU due to runtime constraints.
The assumptions and limitations of our work are as follows: we assume a simple energy model, specifically the one studied in Ref. [28]. We emphasise that our work is a benchmarking study, with the graphene defect problem serving as a representative test-case. For a more detailed discussion of the underlying use-case, its applicability to real-world problems, and comparisons between quantum annealing and density functional theory: we direct the reader to Ref. [28]. In our work, the problem constraint is encoded into the QUBO formalism using penalty terms, rather than solving a hard-constrained version of the problem. Furthermore, our study was conducted using some of the best commercially available quantum devices at the time of experimentation, which took place in January 2025. As of January 2026 the quantum devices used in this investigation are still accessible. Ibm_fez using the Heron r2 processor still has the joint highest number of qubits of any IBM device. Newer Heron r3 devices have lower two qubit error rates but these error rates are of the same order of magnitude. Readout errors in the Heron r3 devices are of an order of magnitude lower than those of ibm_fez. IBM recently announced the Nighthawk r1 processor with square lattice connectivity instead of heavy hex. This change would reduce SWAP overheads and enable shorter VQE circuits for fully connected problems such as those considered here. D-Wave’s Advantage 6.4 system remains accessible, while newer Advantage2 devices are now available with improved qubit connectivity. Nevertheless, our work remains relevant due to the generality of our benchmarking framework and the algorithms considered. In particular, the algorithmic bottlenecks identified persist despite hardware upgrades. We did not apply error mitigation or other techniques that could improve algorithm performance (as listed in Section VI).
In Section II we present other recent efforts in benchmarking optimisation algorithms and real world implementations. We then formulate our specific problem in Section III. In Section IV we describe our methodology for the cross-platform benchmarking including the chosen metrics, whose results we present in Section V. Finally, in Section VI we discuss open questions and future directions. Additional details of our methodology are included in the Appendices.
II Related Work
@@
More specifically, Ref. [28] demonstrates how QAA can be applied to solve a version of the graphene defect use-case formulated as a Densest k-Subgraph (DkS) problem (see Appendix A for more detail on its formulation and complexity). This problem is NP-hard for general and bipartite graphs of degree up to [67]. In [28], an atom lattice is encoded as a fully-connected QUBO and executed on quantum annealers. In our work we execute this on larger lattice sizes, using quantum annealers as well as gate-based quantum computers and classical methods. Another related piece of work [69] studies a slightly different formulation of the problem where the goal is to minimise the number of dangling bonds (i.e. connections between vacancies and carbon atoms in the lattice), again executed on a quantum annealer. The same DkS problem as in [28] has also been investigated in a biology-inspired test-case [15] on randomised bipartite graphs, where the topology of the D-Wave device is identified as the main limitation for applying it on quantum hardware.
@@
Ref. [25] provides a comprehensive review of the current state of quantum optimisation, highlighting the challenges and future prospects of the field. The review outlines best practices for making fair comparisons when solving optimisation problems with different methods. The authors showcase recent publications where optimisation problems with varying numbers of qubits and problem connectivity (known as density) are solved using VQAs, comparing each solution’s quality through the approximation ratio metric. The most common problem connectivity involves 3-regular graphs (where each node is connected to three others), which aligns well with quantum hardware layouts. Studies that explore higher qubit counts typically maintain this low problem connectivity. Increasing connectivity for larger qubit systems requires additional SWAP gates, leading to increased hardware noise. However, Refs. [5, 48, 88, 37] all consider Sherrington-Kirkpatrick problems with 100% density, which can be reformulated as QUBO problems. These studies emphasize both the challenge and the importance of choosing standardised metrics to ensure that benchmarks are fair and results are comparable. The chosen set of metrics must fully capture the diverse behaviors of different algorithms. Our work applies the best practices outlined in this review to a particular problem, which stands out in terms of qubit count and high density compared to the other highlighted studies. We achieve a competitive mean approximation ratio with VQE on a QPU relative to the results reported in Table IV of Ref. [25]. In terms of problem size, comparison with Table IV shows that our work is competitive with the state of the art in achieving the largest qubit counts for problems solved at 100% density. While Ref. [25] primarily focuses on gate-based algorithms, we also address questions regarding how QAAs can be fairly compared to variational gate-based methods.
@@
Other studies [38, 46] also compare QAAs and gate-based variational methods for specific use-cases. However, a comprehensive study that includes time comparisons, solution quality, and scaling analysis has yet to be conducted. Ref. [51] compares the Quantum Approximate Optimisation Algorithm (QAOA) to quantum annealing and claims that QAOA, when run on a real gate-based device, can outperform quantum annealing on a 5-regular max-cut graph problem. On the other hand, Ref. [14] argues that these comparisons could be seen as ambiguous, explaining that the reported results from QAOA were enhanced by post-selection. These studies further highlight the need for transparent and fair benchmarking. In our work, we expose post-selection impacts - both in our metrics and distributions - as well as hyperparameter tuning, embedding overheads, and runtime break downs in a transparent and reproducible manner.
Recently, Ref. [92] benchmarks gate-based hardware using the Linear Ramp QAOA (LR-QAOA) method [93]. The advantage of using LR-QAOA over traditional near-term gate-based algorithms is that it does not involve variational parameters. The difficulty of optimising variational parameters, especially on quantum hardware, is demonstrated in our work.
While prior work typically focuses on sparse problems on gate-based devices, we solve dense QUBOs with up to 72 variables. We demonstrate best practices in our cross-platform benchmarking, enabling the evaluation of a gate-based variational algorithm and quantum annealing within a unified framework.
III Problem Formulation
@@
We use the problem of finding the minimum energy of defective graphene structures as an illustrative example of our benchmarking strategy. This problem is discussed in more detail in Appendix A. We model the system as an -site hexagonal structure, as shown visually in Figure 1 ( in this example). The first graphene structure explored is a supercell, a unit cell containing two carbon atoms repeated in a arrangement, which corresponds to an variable QUBO. We also investigate other supercells, which translate to QUBOs of varying size. Each (binary) QUBO variable represents a site containing either a carbon atom or a vacancy.
@@
Each of the carbon atoms in Figure 1 can be removed, creating a vacancy that results in the breaking of carbon-carbon bonds. The breaking of these bonds increases the energy of the system’s ground state configuration. The energy increase depends on where the vacancies are introduced. In particular, and following Ref. [28], the energy of the system depends only on the number of remaining carbon-carbon bonds. In this model we impose periodic boundary conditions. Referring to 1, the periodic boundary conditions can be visualised as connecting atoms on the opposite side of the cell. As a result, the lattice can be represented as a 3-regular graph. The supercell’s 2D geometry is maintained for all configurations.
@@
IV Methodology for Multi-Algorithm Benchmarking
@@ IV.1 Algorithms and Parameters
Many of the methods that we employ here involve hyperparameters whose values must be set. Our hyperparameter search finds the set of parameters for each method that return the highest probability of sampling the ground state. A hyperparameter search could be performed aiming to optimise performance with respect to another metric such as runtime. Our hyperparameter optimisation is a grid-based search, conducted by defining feasible search spaces and checking different combinations of values. The results of the hyperparameter search for all methods in this section are provided in Appendix B. Figures 6, 7, 8, and 9 illustrate the robustness of our hyperparameter search: they display not only the optimal parameter settings but also how performance varies across the full set of tested parameters, highlighting the sensitivity of the methods to parameter choices.
@@ IV.2 Post-Selection
To perform post-selection for our use-case with a single constraint, a constant number of bitstrings must be checked, equal to the number of samples (shots), which have entries each. Therefore, the scaling of this method is linear in . The time required for post-selection at the QUBO problem size, was found to be approximately for samples, which is the maximum number of samples produced as output in any of our experiments. However, for larger problems, the number of samples may need to scale with to achieve non-zero probability of the optimal solution. As increases, this probability decreases due to increased noise from using more qubits, which we have confirmed experimentally.
@@
IV.3 Performance Metrics
Defining good benchmarking strategies for measuring performance of optimisation algorithms has been studied in both quantum and classical computing contexts [104]. In general, these strategies fall under two distinct categories. The first category involves allocating a fixed amount of resources, such as time or energy, and measuring the quality of solution obtained, using a well-defined metric for it. The second involves measuring the amount of a resource, using an appropriate metric, that is necessary to acquire a solution that achieves a pre-specified threshold (or target) for the quality of solution.
@@ IV.3 Performance Metrics
In our case we calculate AR using the solutions after post-selection. The reason we do this is because before post-selection we can have energies outside the range . These and values are calculated using constrained brute-force techniques. The AR places the average energy found within the context of the whole energy landscape of the problem. We note that there is some variation in the name and precise definition used in the literature for this metric. When exploring new large problems, the theoretical minimum and maximum values of energy may not be readily available without having to solve using an exact method. An option for larger problems is to use, in place of the AR, the average energy and the minimum energy obtained experimentally by a given method.
@@
Having discussed the selected metrics, we also note a number of other metrics that are used in the wider literature. Ref. [46], considers the number of violated constraints. For our QUBO there is only one constraint specified - the number of vacancies. The number of times it is violated could be reflected in a ‘validity’ metric: if many solutions violate the constraint then the validity returned will be low. We instead choose to encode information about constraint violation by post-selecting the data to remove solutions that do not obey the constraint, as described above. A metric specific to the quality of the quantum annealing process in current devices is the percentage of broken chains. We report this value in Appendix E, Figure 18d.
Another metric often used is the time-to-solution (TTS), where this is the time taken for a method to find the optimal solution once with a desired accuracy (often ). TTS, as defined in Ref. [44, 56], is:
@@
Here we create our collection of solutions by aggregating shots across many experiments. This means that the total number of solutions for quantum annealing and VQE is equal to , with being the number of times the experiment is repeated and the number of shots per experiment. While in simulated annealing and random sampling it is equal to the number of times we run the algorithm. For the definitions of shots and experiments see Section IV.1. This is necessary with quantum annealing as the D-Wave devices have a maximum QPU access time, and in order to increase the number of shots to a sufficiently large number, one needs to run the experiment again[82].
@@
V Results
We first solve the variable QUBO and analyse our methods performance using the metrics defined in Section IV.3. Figure 3 shows the energy probability distributions obtained by each method before post-selection.